# A Hierarchical Approach to Quantum Many-Body Systems in Structured Environments

### v1

@@ III.1 Driven-dissipative many-emitter systems
III.2 Many-body electronic systems coupled to non-Markovian optical environments
@@
The electronic structure of most organic molecules is dominated by covalent bonds or conjugated $\pi$-systems, requiring many-body methods to faithfully model their electronic properties. An excitation in such a system cannot be represented by the excitation events of a single electron. Instead, the surrounding electronic cloud reacts and screens the polarization induced by the external field – every optical excitation in a molecule or solid is inherently a many-body process. Electronic correlation can increase to bizarre levels for systems including transition metals, which contribute strongly localized d and f-orbitals shielded from molecular orbitals. Moving two electrons to such a state costs a considerable amount of energy, as they strongly repel each other, resulting in electronic motion that is strongly correlated. A prototypical approach to describe such a system is given by the Hubbard model
@@
Naturally, the electronic structure is embedded in an environment of optical (or vibro-phononic) modes, resulting in the emergence of polaritonic states if coupled sufficiently strong. Our BBGKY-HEOM approach allows to consistently describe and monitor how the dynamic of the lossy cavity mode imprints correlations in the electronic many-body system and vice versa. The corresponding minimal Hamiltonian for such a Cavity-Fermi-Hubbard system [68, 69]
$\begin{split}\hat{H}&=\hat{H}_{H}+\hbar\omega(\hat{a}^{\dagger}\hat{a}+\frac{1}{2})\\ &-\sqrt{\frac{\hbar\omega}{2\varepsilon_{0}V}}(\hat{a}+\hat{a}^{\dagger})\textbf{e}_{cav}\cdot\hat{\boldsymbol{\mu}}+\frac{1}{2\varepsilon_{0}V}(\textbf{e}_{cav}\cdot\hat{\boldsymbol{\mu}})^{2},\end{split}$ (18)
includes self-polarization contributions of the localized dipole moment $\hat{\boldsymbol{\mu}}=\textbf{e}_{chain}\sum_{\sigma}\sum_{i=1}^{N}r_{i}\hat{c}_{i,\sigma}^{\dagger}\hat{c}_{i,\sigma}$. The local positions $r_{i}\in\{-\frac{N-1}{2}\Delta,...,0,...,+\frac{N-1}{2}\Delta\}$ as well as their spacing $\Delta$ are defined with respect to the center of the PZW gauge [42, 43]. We will use $J=1$, $\omega=0.5$, assign a cavity loss of $\kappa=0.2$, a coupling strength of $g=\sqrt{\frac{\hbar\omega}{2\varepsilon_{0}V}}=0.1$, and disregard the self-polarization for the moment.
@@
Charge will not only localize once but randomly tunnel in and out of the impurity. The dynamic originating from the random localization of charge on the dipole moment is shown in Fig. 2 (right) with the corresponding changes in occupation (inset) and time-dependent (gray dotted) potential. Clearly, hitting a resonance condition such as at $Jt=40$ can result in large deviations from the equilibrium state without noticeable influence on the quality of the BBGKY-HEOM predictions. One strength of the BBGKY-HEOM is therefore to resolve the many-body electronic dynamic that originates from time-dependent modulations which allows us to simulate e.g. charge noise, dynamic screening, or triplet-triplet annihilation explicitly – a feature that we plan to leverage further in the future.
@@
Figure 3: Strong dynamic in Cavity-Fermi-Hubbard model: Occupation for second and last site, both initially doubly occupied, over time. Sites 0 and 2 are initially empty, resulting in rapid charge-oscillations between the sites and a quick build-up of 3-body correlations. The BBGKY-HEOM becomes increasingly worse with ongoing time which is accompanied with the appearance of negative eigenvalues in the 1RDM that could potentially lead to instabilities.
@@
Let us end this section by illustrating the limitations of the BBGKY hierarchy. The 4 sites of our Hubbard-chain host again 4 electrons but we force them initially to occupy the 1st and 3rd site only, i.e., our initial state has two doubly occupied and two empty sites. Naturally, this state is far from the correlated ground-state and results in a violent dynamic of density sloshing forward and backward. Fig. 3 illustrates the onsite occupation, i.e., the electronic density, of the initially occupied sites over time. The exact reference calculations (black dashed) follow our predictions closely up to $Jt\approx 20$ before the quality of the BBGKY-HEOM prediction deteriorates. We can relate this sudden loss of accuracy to the appearance of artificial negative eigenvalues in the reduced density matrix (see inset) – a failure caused by the fact that the BBGKY hierarchy is not contraction consistent by default. Contraction consistency refers to the fact that the 2RDM must be obtained from the trace of the approximated 3RDM. Neglecting this condition results in a build-up of negative eigenvalues that will inevitably lead to numerical instabilities. Fortunately, such effects can be circumvented by enforcing contraction consistency via more complex reconstruction functionals $\widetilde{F}_{123}$ as shown in [70, 50]. BBGKY-HEOM is remarkably stable when compared to the bare dynamic of the electronic system alone. We attribute this observation to the fact that the decoherence imprinted by the structured bath penalizes instabilities as large uncontrolled oscillations couple efficiently to the optical mode which dissipates a notable fraction of the energy and thus stabilized the dynamic of the system.
III.3 Many-body dynamic embedded into organic crystals
@@
Many-body electronic systems, such as molecules or (an)organic frameworks, are often embedded in a liquid or crystalline environment. For example, organic molecules embedded in organic crystals can serve as highly coherent emitter[1], quantum memory[71], or organic light-emitting diode. Their environment is a combination of local vibrational, extended phononic, and photonic modes, giving rise to a highly structured and complex spectral function. BBGKY-HEOM provides now the means to retain the complexity of the electronic structure and combine it with the full system-bath dynamic, i.e., extending beyond previous approaches that relied on a simplified description of the electronic system[72].
Fig. 4 demonstrates the impact of the non-Markovian phonon-bath on the dynamic of the dipole moment of the same 4-site Cavity-Fermi-Hubbard model ($U=0.1$) subsequent to a quench induced by a charge fluctuation $V_{i}(0^{+})=\frac{4}{5(i+1)}$. As the nuclei move closer together (further apart) their effective hopping rate increases (decreases), which we model by coupling the phonon-bath to the hopping operator:
@@ IV Conclusion
No quantum system is perfectly isolated and the description of its environment can be essential for our physical understanding. Confined optical or plasmonic modes and vibrational or phononic modes might even challenge the dynamic of a subsystem of interest by strongly coupling to the internal degrees of freedom. Energy deposited in the environment will then not simply decohere but act back on the system, resulting in non-Markovian, or memory dependent, system-bath interaction. The dynamic of the system has often been simplified to the degree that the many-body character inherited by every molecule or solid-state impurity has been lost.
In this study, we presented a rigorous hierarchical approach based on the combination of the reduced density-matrix (BBGKY) and system-bath (HEOM) hierarchy. BBGKY-HEOM is flexible enough to tackle typical quantum optical and condensed matter problems, striking a balance between high accuracy and good performance. In particular, we demonstrated that correlated (superradiant) emission and spin squeezing, i.e., dynamic that extends beyond the capabilities of meanfield theory, is accurately captured. We then leveraged the ability to describe many-body electronic systems, here in the form of the Fermi-Hubbard model, to illustrate the interplay between correlated system dynamic and structured baths by simulating different forms of charge noise. We observed again excellent performance for as long as the onsite repulsion U and the electronic dynamic remain moderate. Finally, we demonstrated the major strength of the BBGKY-HEOM by simulating the interplay between the vibro-phononic bath of an organic crystal coupled to the many-body dynamic of the Cavity-Fermi-Hubbard model. Both contributions, i.e., non-Markovian system-bath coupling and many-body dynamics are equally important – clearly stressing the need for a more holistic perspective on solid-state quantum emitters.
@@ Acknowledgements.
We thank Walter Strunz and Göran Johansson for insightful discussions. C.S. acknowledges support from the Swedish Research Council through Grant No. 2016-06059 and funding from the Horizon Europe research and innovation program of the European Union under the Marie Skłodowska-Curie grant agreement no. 101065117. Partially funded by the European Union. Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or REA. Neither the European Union nor the granting authority can be held responsible for them.
@@ Appendix A Spin squeezing
Appendix B Derivation of the BBGKY-HEOM equations
In the following we detail the derivation of the BBGKY-HEOM equation. For the sake of readability we focus on the case of a scalar index, corresponding to the case where the environment is a damped cavity mode. The extension to a vector index is trivial. Starting from the HEOM equation (9) with the auxiliary density matrix
$\rho^{(m,p)}=(ig)^{m}(-ig)^{p}\operatorname{Tr}_{C}\left((a^{\dagger})^{p}a^{m}\rho_{tot}\right),$ (20)
where $\rho_{tot}$ is the total state of system and cavity mode. We now want to employ the BBGKY hierarchy to reduce the dimension of the system.
We use the same normalization for the reduced density matrices as in the known BBGKY hierarchy,
$\begin{split}F_{123}=&N(N-1)(N-2)\operatorname{Tr}_{C}(\rho_{123C}),\\ F_{12}=&N(N-1)\operatorname{Tr}_{C}(\rho_{12C}),\\ F_{1}=&N\operatorname{Tr}_{C}(\rho_{1C}).\end{split}$ (21)
Following Eq. (20) we define the auxiliary matrices as
$F_{1..N}^{(m,p)}=(ig)^{m}(-ig)^{p}\operatorname{Tr}_{C}\left((a^{\dagger})^{p}a^{m}F_{1..N,C}\right).$ (22)
@@ Appendix B Derivation of the BBGKY-HEOM equations
Their evolution equation can be derived from the HEOM Eq. (9) by tracing over all but $N$ particles and assuming that the density matrix is invariant under particle exchange:
@@
$\begin{split}\dot{F}_{12}^{(m,p)}=&\partial_{t}N(N-1)(ig)^{n}(-ig)^{m}\operatorname{Tr}_{3...N,C}\left((a^{\dagger})^{m}a^{n}\rho\right),\\ =&N(N-1)\operatorname{Tr}_{3..N}(\dot{\rho}^{(n,m)}),\\ =&-i[H_{1}+H_{2},F_{12}^{(n,m)}]\\ &-\left[(n-m)i\Delta+(m+n)\kappa\right]F_{12}^{(n,m)}\\ &\quad\!+g^{2}\Big{(}n(L_{1}+L_{2})F_{12}^{(n-1,m)}+\operatorname{Tr}_{3}(L_{3}F_{123}^{(m-1,p)})\\ &\quad\!\!m\rho^{(n,m-1)}(L_{1}^{\dagger}+L_{2}^{\dagger})+\operatorname{Tr}_{3}(L_{3}^{\dagger}F_{123}^{(m,p-1)})\Big{)}\\ &\quad\!+\left[F_{12}^{(n+1,m)},L_{1}^{\dagger}+L_{2}^{\dagger}\right]+\left[L_{1}+L_{2},F_{12}^{(n,m+1)}\right].\end{split}$ (23)
Above we used that
$\begin{split}\operatorname{Tr}_{3...N}([\sum_{i=1}^{N}H_{i},\rho]=&[H_{1}+H_{2},\rho_{12}]\\ &+\sum_{i=3}^{N}\operatorname{Tr}_{3...N}(H_{i}\rho)-\operatorname{Tr}_{3...N}(\rho H_{i}),\\ =&[H_{1}+H_{2},\rho_{12}]\\ &+\sum_{i=3}^{N}\operatorname{Tr}_{3...N}(H_{i}\rho)-\operatorname{Tr}_{3...N}(H_{i}\rho),\\ =&[H_{1}+H_{2},\rho_{12}].\end{split}$ (24)
@@
In the next step we use a cluster expansion to express the term $F_{123}^{(m,p)}$ in terms of the reduced density one- and two-body matrices. The key approximation that enables this is the neglect of higher order correlations. The cluster expansion amounts to expressing the 3-body density matrix including the cavity as:
$\begin{split}\rho_{123C}=&\rho_{1}\otimes\rho_{2}\otimes\rho_{3}\otimes\rho_{C}\\ &+\rho_{1}\otimes\rho_{2}\otimes\Gamma_{3C}+\rho_{1}\otimes\Gamma_{23}\otimes\rho_{C}+\Gamma_{12}\otimes\rho_{3}\otimes\rho_{C}\\ &+\Gamma_{13}\otimes\rho_{2}\otimes\rho_{C}+\Gamma_{1C}\otimes\rho_{2}\otimes\rho_{3}+\rho_{1}\otimes\Gamma_{2C}\otimes\rho_{3}\\ &+\Gamma_{12}\otimes\Gamma_{3C}+\Gamma_{1C}\otimes\Gamma_{23}+\Gamma_{13}\otimes\Gamma_{2C}\\ &+\rho_{1}\otimes\Gamma_{23C}+\Gamma_{13C}\otimes\rho_{2}+\Gamma_{12C}\otimes\rho_{3}+\Gamma_{123}\otimes\rho_{C}\\ &+\Gamma_{1234C}\end{split}$ (25)
@@
We now neglect 3-body correlations within the system $\Gamma_{123}=0$ as well as four body correlations $\Gamma_{123C}=0$ to obtain
@@
$\begin{split}\widetilde{F}^{({m},{n})}_{123}&=4\frac{(N-1)(N-2)}{N^{3}}\operatorname{Tr}(F_{1}^{({m},{n})})F_{1}F_{2}F_{3}\\ &+\frac{N-2}{N}\left(F_{12}F_{3}^{({m},{n})}+F_{2}^{({m},{n})}F_{13}+F_{1}^{({m},{n})}F_{23}\right)\\ &+\frac{N-2}{N}\left(F_{1}F_{23}^{({m},{n})}+F_{13}^{({m},{n})}F_{2}+F_{12}^{({m},{n})}F_{3}\right)\\ &-\frac{N-2}{N^{2}}\operatorname{Tr}(F_{1}^{({m},{n})})\left(F_{12}F_{3}+F_{13}F_{2}+F_{1}F_{23}\right)\\ &-2\frac{(N-2)(N-1)}{N^{2}}\Big{(}F_{1}^{({m},{n})}F_{2}F_{3}\\ &\quad+F_{1}F_{2}^{({m},{n})}F_{3}+F_{1}+F_{2}F_{3}^{({m},{n})}\Big{)}.\end{split}$ (26)

### v2

@@ III.2 Few-emitter lasing
III.3 Many-body electronic systems coupled to non-Markovian optical environments
Figure 4: Fermi-Hubbard model coupled to lossy cavity mode: (Right) Dipole moment driven by time-dependent potential $V_{i}(t)=\frac{f(t)}{5(i+1)}$, where f(t) is a step function as shown by the grey dashed line and $U=0.1$. The example simulates charge noise due to the movement of electrons in the vicinity of the molecule. Inset illustrates onsite occupation.
@@ III.3 Many-body electronic systems coupled to non-Markovian optical environments
We will now relax the simplification of the emitter and instead combine the dynamics of a correlated electronic system with a structured system-bath coupling. The electronic structure of most organic molecules is dominated by covalent bonds or conjugated $\pi$-systems, requiring many-body methods to faithfully model their electronic properties. An excitation in such a system cannot be represented by the excitation events of a single electron. Instead, the surrounding electronic cloud reacts and screens the polarization induced by the external field – every optical excitation in a molecule or solid is inherently a many-body process. Electronic correlation can reach substantial levels for e.g. transition metals, which contribute strongly localized $d$ and $f$ orbitals shielded from molecular orbitals. Moving two electrons to such a state costs a considerable amount of energy, as they strongly repel each other, resulting in electronic motion that is strongly correlated. A prototypical approach to describe such a system is given by the Hubbard model
@@
Naturally, the electronic structure is embedded in an environment of optical (or vibro-phononic) modes, resulting in the emergence of polaritonic states if coupled sufficiently strong. Our BBGKY-HEOM approach allows to consistently describe and monitor how the dynamics of the lossy cavity mode imprints correlations in the electronic many-body system and vice versa. The corresponding minimal Hamiltonian for such a Cavity-Fermi-Hubbard system [89, 90]
$\begin{split}\hat{H}&=\hat{H}_{H}+\hbar\omega(\hat{a}^{\dagger}\hat{a}+\frac{1}{2})\\ &-\sqrt{\frac{\hbar\omega}{2\varepsilon_{0}V}}(\hat{a}+\hat{a}^{\dagger})\textbf{e}_{cav}\cdot\hat{\bm{\mu}}+\frac{1}{2\varepsilon_{0}V}(\textbf{e}_{cav}\cdot\hat{\bm{\mu}})^{2},\end{split}$ (24)
includes self-polarization contributions of the localized dipole moment $\hat{\bm{\mu}}=\textbf{e}_{chain}\sum_{\sigma}\sum_{i=1}^{N}r_{i}\hat{c}_{i,\sigma}^{\dagger}\hat{c}_{i,\sigma}$. The local positions $r_{i}\in\{-\frac{N-1}{2}\Delta,...,0,...,+\frac{N-1}{2}\Delta\}$ as well as their spacing $\Delta$ are defined with respect to the center of the Power-Zienau-Wooley (PZW) gauge [68, 69]. We will use $J=1$, $\omega=0.5$, assign a cavity loss of $\kappa=0.2$, a coupling strength of $g=\sqrt{\frac{\hbar\omega}{2\varepsilon_{0}V}}=0.1$, and disregard the self-polarization for the moment. We emphasize that extensions to ultra-strong coupling [91] require the careful consideration of self-polarization terms [68, 69].
@@
Charge will not only localize once but randomly tunnel in and out of the impurity. The dynamics originating from the random localization of charge on the dipole moment is shown in Fig. 4 (right) with the corresponding changes in occupation (inset) and time-dependent (gray dotted) potential. Clearly, hitting a resonance condition such as at $Jt=40$ can result in large deviations from the equilibrium state without noticeable influence on the quality of the BBGKY-HEOM predictions. One strength of the BBGKY-HEOM is therefore to resolve the many-body electronic dynamics that originates from time-dependent modulations which allows us to simulate e.g. charge noise, dynamics screening, or triplet-triplet annihilation explicitly – a feature that we plan to leverage further in the future.
Let us end this section by illustrating the limitations of the BBGKY hierarchy. The 4 sites of our Hubbard-chain host again 4 electrons but we force them initially to occupy the 1st and 3rd site only, i.e., our initial state has two doubly occupied and two empty sites. Naturally, this state is far from the correlated ground-state and results in a violent dynamics of density sloshing forward and backward. Fig. 5 illustrates the onsite occupation, i.e., the electronic density, of the initially occupied sites over time.
@@
Figure 5: Strong dynamics in Cavity-Fermi-Hubbard model: Occupation for second and last site, both initially doubly occupied, over time. Sites 0 and 2 are initially empty, resulting in rapid charge-oscillations between the sites and a quick build-up of 3-body correlations. The BBGKY-HEOM becomes increasingly worse with ongoing time which is accompanied with the appearance of negative eigenvalues in the 1RDM that could potentially lead to instabilities.
@@
The exact reference calculations (black dashed) follow our predictions closely up to $Jt\approx 20$ before the quality of the BBGKY-HEOM prediction deteriorates. We can relate this sudden loss of accuracy to the appearance of artificial negative eigenvalues in the reduced density matrix (see inset) – a failure caused by the fact that the BBGKY hierarchy is not contraction consistent by default. Contraction consistency refers to the fact that the 2RDM must be obtained from the trace of the approximated 3RDM. Neglecting this condition results in a build-up of negative eigenvalues that will inevitably lead to numerical instabilities. Fortunately, such effects can be circumvented by enforcing contraction consistency via more complex reconstruction functionals $\widetilde{F}_{123}$ as shown in [92, 66]. BBGKY-HEOM is remarkably stable when compared to the bare dynamics of the electronic system alone. We attribute this observation to the fact that the decoherence imprinted by the structured bath penalizes instabilities as large uncontrolled oscillations couple efficiently to the optical mode which dissipates a notable fraction of the energy and thus stabilized the dynamics of the system.
III.4 Many-body dynamics embedded into organic crystals
@@
Many-body electronic systems, such as molecules or (an)organic frameworks, are often embedded in a liquid or crystalline environment. For example, organic molecules embedded in organic crystals can serve as a highly coherent emitter [1], quantum memory [93], or organic light-emitting diode. Their environment is a combination of local vibrational, extended phononic, and photonic modes, giving rise to a highly structured and complex spectral function. BBGKY-HEOM provides now the means to retain the complexity of the electronic structure and combine it with the full system-bath dynamic, i.e., extending beyond previous approaches that relied on a simplified description of the electronic system [94].
Fig. 6 demonstrates the impact of the non-Markovian phonon-bath on the dynamics of the dipole moment of the same 4-site Cavity-Fermi-Hubbard model ($U=0.1$) subsequent to a quench induced by a charge fluctuation $V_{i}(0^{+})=\frac{4}{5(i+1)}$. As the nuclei move closer together (further apart) their effective hopping rate increases (decreases), which we model by coupling the phonon-bath to the hopping operator:
@@ IV Conclusion
No quantum system is perfectly isolated, and the description of its environment can be essential for our physical understanding. Confined optical or plasmonic modes and vibrational or phononic modes might even challenge the dynamics of a subsystem of interest by strongly coupling to the internal degrees of freedom. Energy deposited in the environment will then not simply decohere but act back on the system, resulting in non-Markovian, or memory dependent, system-bath interaction. The dynamics of the system has often been simplified to the degree that the many-body character inherited by every molecule or solid-state impurity has been lost.
In this study, we presented a rigorous hierarchical approach based on stacking of the reduced density-matrix (BBGKY) and system-bath (HEOM) hierarchies together. BBGKY-HEOM is flexible enough to tackle critical quantum optical and condensed matter problems, striking a balance between high accuracy and good performance. In particular, we demonstrated that correlated (superradiant) emission and spin squeezing, i.e., dynamics that extend beyond the capabilities of mean-field theory, are accurately captured.
Bolstered by this excellent performance, we move to describe few-emitter lasing of methylene-blue molecules in plasmonic nanocavities. Explicit treatment of the structured vibrational bath uncovers resonant enhancements in the lasing process that go beyond the paradigmatic incoherent drive model. To the best of our knowledge, this effect has not been discussed before. BBGKY-HEOM thus demonstrates its potential for a more detailed understanding of light-matter interaction in general, and few-emitter lasing in particular.
Many-body electronic systems, here in the form of the Fermi-Hubbard model, are critical in modern condensed matter theory. We illustrate the interplay between correlated system dynamics and structured baths by simulating different forms of charge noise. Again, excellent performance is observed as long as the onsite repulsion $U$ and the electronic dynamic remain moderate. Finally, we demonstrated the major strength of BBGKY-HEOM by simulating the interplay between the vibro-phononic bath of an organic crystal coupled to the many-body dynamics of the Cavity-Fermi-Hubbard model. Both contributions, i.e., non-Markovian system-bath coupling and many-body dynamics are equally important – clearly stressing the need for a more holistic perspective on solid-state quantum emitters.
Our hierarchical approach can be rigorously informed from first principles – bridging previously disconnected research fields. In this way, BBGKY-HEOM provides a powerful tool that can be used alongside established methods (like cumulant expansions, density-functional theory, etc.) to gather a more detailed and holistic understanding. Especially promising examples include molecules interacting with optical fields [95], may it be for quantum technological applications [1, 96], polariton condensation [97], or few-emitter lasing [79].
@@ Acknowledgements.
We thank Walter Strunz, Alexander Eisfeld, and Göran Johansson for insightful discussions. C.S. acknowledges support from the Swedish Research Council through Grant No. 2016-06059 and funding from the Horizon Europe research and innovation program of the European Union under the Marie Skłodowska-Curie grant agreement no. 101065117. Partially funded by the European Union. Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or REA. Neither the European Union nor the granting authority can be held responsible for them.
@@ Appendix A Spin squeezing
Appendix B Derivation of the HEOM
We sketch here a derivation of the HEOM equations. We assume that the system is coupled to the environment via operator $q$, such that in an interaction picture with respect to the free Hamiltonian of the environment we have
$\begin{split}H_{tot}=&H_{sys}+q\left(B(t)+B^{\dagger}(t)\right),\\ B(t)=&\sum_{\lambda}g_{\lambda}e^{-i\omega_{\lambda}t}a_{\lambda}.\end{split}$ (26)
The response of the system to the environment is given by the bath correlation function $\alpha(t-s)=\langle B(t)B^{\dagger}(s)\rangle$. For HEOM we need that the bath correlation function is of the form
$\alpha(t)=\sum_{j=1}^{N_{exp}}G_{j}e^{-W_{j}t},$ (27)
where $G_{j}$ and $W_{j}$ may be complex numbers. For increased readability we perform the next steps for $N_{exp}=1$, but the extension to multiple exponentials is straight forward. The reduced dynamics of the system can be written in the path integral form
We define auxiliary states
$\rho^{(n,m)}(q,q^{\prime},t)=$ $\int\mathrm{d}q_{0}\int\mathrm{d}q_{0}^{\prime}\int\mathcal{D}q\int\mathcal{D}q^{\prime}\,$
$\dot{\rho}^{(n,m)}(t)=$ $-i[H,\rho^{(n,m)}]-(nW+mW^{*})\rho^{(n,m)}$
$+[\rho^{(n+1,m)},q]+[q,\rho^{(n,m+1)}]$
$+nGq\rho_{t}^{(n-1,m)}+mG^{*}\rho^{(n,m-1)}q.$ (31)
The equations presented in the main text are a generalization of the above equations in two respects. First, we allow for multiple exponential terms in the expansion for the bath correlation function, and second we allow for the case $q\neq q^{\dagger}$ [48]. Additionally, it can be shown that
$\rho^{(n,m)}=i^{n}(-i)^{m}\operatorname{Tr}_{env}\left((B^{\dagger}(t))^{m}B(t)^{n}\rho_{tot}\right).$ (32)
Appendix C Derivation of the BBGKY-HEOM equations
In the following we detail the derivation of the BBGKY-HEOM equation. For the sake of readability we again focus on the case of a scalar index, corresponding to a single (global) bath that all particles couple to and a single exponential in the bath correlation function. The situation where each particle in the many-body system couples to its own (local) bath is explained in Appendix D. The extension to a vector index is straight forward. We start from the HEOM equation (II.2) and want to employ the BBGKY hierarchy to reduce the dimension of the system.
Following (11) we define the reduced 2-body auxiliary matrices as
$F_{12}^{(n,m)}=N(N-1)\operatorname{Tr}_{3..N}(\rho^{(n,m)}).$ (33)
@@ Appendix C Derivation of the BBGKY-HEOM equations
Their evolution equation can be derived from the HEOM Eq. (II.2) by tracing over all but $2$ particles and assuming that the density matrix is invariant under particle exchange:
@@
$\begin{split}\dot{F}_{12}^{(n,m)}=&N(N-1)\operatorname{Tr}_{3..N}(\dot{\rho}^{(n,m)}),\\ =&-i[H_{1}+H_{2},F_{12}^{(n,m)}]\\ &-\left[(n-m)i\Delta+(m+n)\kappa\right]F_{12}^{(n,m)}\\ &\quad\!+Gn\left((L_{1}+L_{2})F_{12}^{(n-1,m)}+\operatorname{Tr}_{3}(L_{3}F_{123}^{(n-1,m)})\right)\\ &\quad\!\!+G^{*}m\left(F_{12}^{(n,m-1)}(L_{1}^{\dagger}+L_{2}^{\dagger})+\operatorname{Tr}_{3}(L_{3}^{\dagger}F_{123}^{(n,m-1)})\right)\Bigg{)}\\ &\quad\!+\left[F_{12}^{(n+1,m)},L_{1}^{\dagger}+L_{2}^{\dagger}\right]+\left[L_{1}+L_{2},F_{12}^{(n,m+1)}\right].\end{split}$ (34)
Above we used that operators which are traced over can be cyclically moved under the trace, such that
$\begin{split}\operatorname{Tr}_{3...N}([\sum_{i=1}^{N}H_{i},\rho]=&[H_{1}+H_{2},\rho_{12}]\\ &+\sum_{i=3}^{N}\operatorname{Tr}_{3...N}(H_{i}\rho)-\operatorname{Tr}_{3...N}(\rho H_{i}),\\ =&[H_{1}+H_{2},\rho_{12}].\end{split}$ (35)
In addition, by utilizing the particle exchange symmetry of $\rho^{(n,m)}$, partial traces over operators can be combined with the help of the particle exchange operator $P_{3j}$ for arbitrary $j\in[4,N]$:
$\begin{split}\operatorname{Tr}_{3..N}(L_{j}\rho^{(m-1,p)})=&\operatorname{Tr}_{3..N}(P_{3j}^{2}L_{j}P_{3j}^{2}\rho^{(m-1,p)}),\\ =&\operatorname{Tr}_{3..N}(L_{3}P_{3j}\rho^{(m-1,p)}P_{3j}),\\ =&\operatorname{Tr}_{3..N}(L_{3}\rho^{(m-1,p)}),\\ =&\frac{\operatorname{Tr}_{3}(L_{3}F_{123}^{(m-1,p)})}{N(N-1)(N-2)}.\end{split}$ (36)
@@
In the next step we use a cluster expansion to express the term $F_{123}^{(n,m)}$ in terms of the reduced density one- and two-body matrices. The key approximation that enables this is the neglect of higher order correlations. Combing Eq. (32) and Eq. (11) we can express the 3-body density matrix as
$\begin{split}F_{123}^{(n,m)}=&N(N-1)(N-2)i^{n}(-i)^{m}\\ &\times\operatorname{Tr}_{4..N,env}\left((B^{\dagger}(t))^{m}B(t)^{n}\rho_{tot}\right).\end{split}$ (37)
The cluster expansion amounts to expressing the 3-body density matrix including the environment as:
$\begin{split}\rho_{123E}=&\rho_{1}\otimes\rho_{2}\otimes\rho_{3}\otimes\rho_{E}\\ &+\rho_{1}\otimes\rho_{2}\otimes\Gamma_{3E}+\rho_{1}\otimes\Gamma_{23}\otimes\rho_{E}+\Gamma_{12}\otimes\rho_{3}\otimes\rho_{E}\\ &+\Gamma_{13}\otimes\rho_{2}\otimes\rho_{E}+\Gamma_{1E}\otimes\rho_{2}\otimes\rho_{3}+\rho_{1}\otimes\Gamma_{2E}\otimes\rho_{3}\\ &+\Gamma_{12}\otimes\Gamma_{3E}+\Gamma_{1E}\otimes\Gamma_{23}+\Gamma_{13}\otimes\Gamma_{2E}\\ &+\rho_{1}\otimes\Gamma_{23E}+\Gamma_{13E}\otimes\rho_{2}+\Gamma_{12E}\otimes\rho_{3}+\Gamma_{123}\otimes\rho_{E}\\ &+\Gamma_{1234E}\end{split}$ (38)
@@
We now neglect 3-body correlations within the system $\Gamma_{123}=0$ as well as four body correlations $\Gamma_{123E}=0$ to obtain
@@
$\begin{split}\widetilde{F}^{({n},{m})}_{123}&=4\frac{(N-1)(N-2)}{N^{3}}\operatorname{Tr}(F_{1}^{({n},{m})})F_{1}F_{2}F_{3}\\ &+\frac{N-2}{N}\left(F_{12}F_{3}^{({n},{m})}+F_{2}^{({n},{m})}F_{13}+F_{1}^{({n},{m})}F_{23}\right)\\ &+\frac{N-2}{N}\left(F_{1}F_{23}^{({n},{m})}+F_{13}^{({n},{m})}F_{2}+F_{12}^{({n},{m})}F_{3}\right)\\ &-\frac{N-2}{N^{2}}\operatorname{Tr}(F_{1}^{({n},{m})})\left(F_{12}F_{3}+F_{13}F_{2}+F_{1}F_{23}\right)\\ &-2\frac{(N-2)(N-1)}{N^{2}}\Big{(}F_{1}^{({n},{m})}F_{2}F_{3}\\ &\quad+F_{1}F_{2}^{({n},{m})}F_{3}+F_{1}+F_{2}F_{3}^{({n},{m})}\Big{)}.\end{split}$ (39)